A capacitor of capacitance C is connected with a resistor of resistance 2r and a battery of emf E and resistance r in a series in a circit.How much heat is inside the battery by the time capacitor is 50percent charged??

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

let voltaage across capacitor=V(t)

V(t)=E(1-e-t/(3RC))

I(t)=CdV(t)/dt

answer=∫(I(t))2r dt, from t=0 to t=T, where V(T)=E/2